Find the equation of a circle with centers (1, 1) and contains points (2, 3)
1
Expert's answer
2021-01-26T11:58:32-0500

Equation of a circle centered at "(1,1)" with radius "R" is "(x-1)^2+(y-1)^2=R^2". Now to find the radius, we can just use the fact that "(2,3)" lies on it :

"(2-1)^2 + (3-1)^2 = 5=R^2"

Therefore the equation is "(x-1)^2+(y-1)^2=5" (or in complex number notation "|z-1-i|=\sqrt{5}").